4. Generic Corp. is a publicly traded company. It has 20 million shares trading at $15/share and its book value of equity is $150 million. The firm also has book value of debt of $100 million and market value of debt of $120 million. The cost of equity for the company is 10%, the pre-tax cost of debt is 4% and the marginal tax rate is 40%. What is the cost of capital? О 6.36% О 7.00% 7.11% О 7.35% None of the above
Answer #1

Market value of Equity = 20* 15 = 300 million
Market value of Debt = 120 million
Total Value = 300+120 = 420

Cost of capital = Equity/Value * Cost of Equity + Debt/Value * Cost of Debt*(1-Tax rate) = 300/420 * 10% + 120/420*4%*(1-40%) = 7.83%

none of these

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well